OpenStudy (anonymous):

Solve y = 3x – 12 by the substitution method. 2x – 3y = -13

OpenStudy (shane_b):

Ok that's pretty simple. The first equation is already in terms of y since it's "y=3x-12". Place (3x-12) any place you see a "y" in the second equation...then solve the second equation for x. Once you know x, place its value back into one of the equations and solve for y.

OpenStudy (shane_b):

We can do it step by step if you like...

OpenStudy (anonymous):

yes please

OpenStudy (mathteacher1729):

Here is a great video on solving systems of linear equations by substitution: http://youtu.be/cwHR_B9zK7k

OpenStudy (shane_b):

Ok. Here are the original equations: y = 3x – 12 2x – 3y = -13 Substitute y in the second equation with 3x-12: 2x – 3(3x-12) = -13 Solve that equation for x to get x=7. Now that you know what x is, place it back into one of the original equations and solve for y: y=3(7)-12 Solving for y you get y = 9.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

thanks!

OpenStudy (shane_b):

yw :)